40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 9 08:23:45 2010, steve, Update, SAFETY, main laser turned off 
    Reply  Thu Sep 9 11:17:15 2010, steve, Update, SAFETY, main laser is back ON P1060813.JPGP1060814.JPGP1060816.JPG
       Reply  Thu Sep 9 16:54:56 2010, rana, Update, SAFETY, main laser is back ON 
Message ID: 3550     Entry time: Thu Sep 9 11:17:15 2010     In reply to: 3549     Reply to this: 3554
Author: steve 
Type: Update 
Category: SAFETY 
Subject: main laser is back ON 


Innolight 2W 1064nm main laser was turned off for more enclosure related work.

The electrician re rooting the power- conduit to the interlock under the floor and bring it through the new concrete tunnel.

Emergency laser power shut off switch was tested at entry door 104M.  It worked fine.


No laser glasses required. Laser will be turned back on around 11 am today.

Atm1,   AC power line to enclosure and interlock  was extended.

Atm2,   24V line to interlock rerouted

Atm3,   job is completed




 The laser is turned on. Injection current set to 0.871A.

Actually some one turned the  laser  off last night and did  not enter it into the elog ! Burned toast award is granted !

Please  come forward voluntarily.

Attachment 1: P1060813.JPG  2.308 MB  | Show | Show all
Attachment 2: P1060814.JPG  2.487 MB  | Show | Show all
Attachment 3: P1060816.JPG  2.263 MB  | Show | Show all
ELOG V3.1.3-